### v5

@@ 5.3 Second Bridge: Quasi-Experimental to Structural
Our merger analysis faces a challenge: the small number of treated markets relative to controls makes it difficult to obtain meaningful and precise parameter estimates. The three mergers, DL-NW, UA-CO, and AA-US, 14 out of 836, 6 out of 774, and 12 out of 779 markets, respectively, each representing a small fraction of the total route network. This limited treatment variation may affect the estimates of efficiency gains and changes in conduct, as well as their precision.
Selecting suitable control markets for merger analysis is a challenging task. The ideal control group should closely mirror the counterfactual evolution of treated markets in the absence of the merger, thereby satisfying the parallel trends assumption that is crucial for DiD estimation. As Ashenfelter et al. (2009) observe, researchers face an inherent tradeoff: control markets must be geographically distant enough to avoid merger spillovers yet similar enough to face comparable demand and cost conditions. This tradeoff presents what Ashenfelter et al. (2009) term the “similarity-independence tension” in control market selection.
Our use of first-difference methods compounds the estimation challenge. Griliches and Mairesse (1999) demonstrate that first-difference estimators, while addressing fixed effects, can substantially reduce identifying variation. Our results confirm this pattern: first-difference specifications consistently yield larger standard errors and less significant coefficients compared to fixed effects approaches (Table C.2). However, fixed effects estimates may be biased due to unobserved market characteristics that are correlated with the merger treatment.
We propose a solution that, in keeping with the overarching goal of the paper, bridges the quasi-experimental literature with structural econometrics. Our approach, which we term synthetic GMM, combines attractive features of synthetic control methods (Abadie and Gardeazabal, 2003) and synthetic DiD (Arkhangelsky et al., 2021) with structural estimation. Like synthetic DiD, our method reweights control markets using pre-treatment price similarity to construct better counterfactuals, thereby strengthening identification. The key innovation is straightforward and tractable: we utilize the market- and time-specific synthetic DiD weights to reweight each market’s contribution to the moment conditions.
@@
3:Construct residualized pre-merger outcome (prices): $\bar{P}_{tr,pre}$ (treated averages), and $P_{co,pre}$ (control matrix)
4:Solve for market weights: $\hat{\omega}=\arg\min_{\omega}\|\bar{P}_{tr,pre}-\omega^{\prime}P_{co,pre}\|_{2}^{2}+\zeta\|\omega\|_{2}^{2}$
5:Solve for time weights: $\hat{\tau}=\arg\min_{\tau}\|\bar{P}_{tr,pre}^{\prime}-\tau^{\top}P_{tr,pre}^{\top}\|_{2}^{2}+\zeta\|\tau\|_{2}^{2}$
6:Define observation weights: $\omega_{mt}=\hat{\omega}_{m}$ if $m\in\mathcal{C}$, else 11; $r_{mt}=\hat{\tau}_{t}$ if pre-treatment, else 11
@@
7:Synthetic moments: $\tilde{g}(\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{\lambda})=\frac{1}{MT}\sum_{i,t}\omega_{mt}\cdot r_{mt}\cdot g(P_{mt},\boldsymbol{W}_{mt},\boldsymbol{Z}_{mt},Q_{mt};(\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{\lambda}))$
8:Synthetic GMM: $(\hat{\boldsymbol{\theta}},\hat{\boldsymbol{\alpha}},\hat{\boldsymbol{\lambda}}):=\arg\min_{(\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{\lambda})}\tilde{g}((\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{\lambda}))^{\top}\hat{W}\tilde{g}((\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{\lambda}))$
9:Bootstrap Inference:
10:for $b=1,\ldots,B$ do
@@
11:  Resample units with replacement (preserving time series and treatment status)
12:  Estimate weights $\hat{\omega}^{(b)},\hat{\tau}^{(b)}$ for bootstrap sample
@@
13:  Estimate synthetic GMM $(\hat{\boldsymbol{\theta}},\hat{\boldsymbol{\alpha}},\hat{\boldsymbol{\lambda}})^{(b)}$ using bootstrap weights
14:end for
@@
15:Compute bootstrap standard errors from $\{(\hat{\boldsymbol{\theta}},\hat{\boldsymbol{\alpha}},\hat{\boldsymbol{\lambda}})^{(b)}\}_{b=1}^{B}$
@@
More formally, our synthetic GMM approach proceeds in two main steps, as shown in the Table 6. First, following Arkhangelsky et al. (2021), we construct market-specific weights $\hat{\omega}$ that align pre-merger price trends between control and treated markets, and time-specific weights $\hat{\tau}$ that balance pre-merger periods. Like synthetic control methods, we find weights such that the weighted average of control market residualized prices closely matches treated market residualized prices: $\sum_{m\in\text{control}}\hat{\omega}_{m}P_{mt}\approx\frac{\sum_{m\in\text{treated}}P_{mt}}{N_{\text{treated}}}$ across all pre-merger periods $t$. The time weights similarly balance temporal variation between the two groups.
@@
Second, we integrate these weights directly into our GMM estimation by reweighting each moment condition, which we refer to as synthetic GMM. Instead of the standard sample moments $\frac{1}{MT}\sum_{m,t}g(\cdot;\cdot)$, we compute synthetic-weighted moments $\frac{1}{MT}\sum_{m,t}\omega_{mt}r_{mt}g(\cdot;\cdot)$, where $\omega_{mt}$ applies market weights to control units and $r_{mt}$ applies time weights to pre-treatment periods. This reweighting ensures that parameter identification emphasizes the most informative control observations while preserving the structural interpretation of synthetic GMM estimates. For inference, we implement block bootstrap resampling that jointly estimates weights and structural parameters, accounting for uncertainty in both components.
@@
Table 7 displays the estimates. The estimates suggest heterogeneity in merger effects across the three airline consolidations, with notably improved precision compared to the estimates in Table 5. The weighting approach yields smaller standard errors across all parameters, while maintaining the same qualitative conclusions about merger impacts. The estimates for the UA-CO merger suggest increased strategic conduct among remaining competitors by 0.049 (p-value 0.1), suggesting enhanced coordination in the post-merger market. Simultaneously, the merger appears to have reduced efficiency, i.e., increased cost, by \$13.5, which represents approximately 5.9% of the average pre-treatment price of \$228.25.
@@
These findings underscore the importance of rigorous merger analysis using a structural approach. The synthetic GMM approach can improve our ability to detect conduct changes that might be missed with standard GMM. The results suggest that airline mergers were not pro-competitive as they generated limited efficiency gains while facilitating coordination.
@@ 6 Conclusion
This paper establishes a connection between quasi-experimental and structural approaches to merger evaluation. We build a two-way methodological bridge. The first bridge shows that the widely used DiD estimation equation emerges as the reduced form of a structural demand and supply equation, which in turn clarifies the reasons why DiD identification may fail in the merger context. The second bridge demonstrates that we can leverage recent advancements in quasi-experimental approaches, particularly synthetic control methods, to enhance the estimation of structural models (synthetic GMM). This bridge reveals that these seemingly distinct methodologies represent different levels of economic detail applied to the same underlying relationships, and offers a unified framework that combines the interpretability of DiD with the robustness of structural identification.
We demonstrate that the DiD coefficient represents a composite parameter conflating cost efficiencies, economies of scale, and strategic conduct changes—effects that can only be separated through structural decomposition. Consequently, mergers present fundamental identification challenges for DiD approaches: when mergers alter strategic behavior, the error structure in DiD specifications becomes endogenous to treatment status, resulting in systematic bias. Our structural framework addresses these limitations while enabling researchers to construct DiD-equivalent aggregate measures from structural parameters, obtaining both methodological robustness and economic interpretability.
@@
Although focused on airlines, our bridging approach offers broad applicability for analyzing mergers in concentrated industries where strategic conduct plays a central role. The methodological framework provides a foundation for robust merger retrospectives that can inform both academic research and antitrust policy (Asker and Nocke, 2021).
@@ B Proofs of the Identification Results
$Q_{mt}$ $=\alpha_{0}+\alpha_{1}P_{mt}+\alpha_{2}^{\prime}\mathbf{Z}_{mt}+\varphi_{mt}\quad\text{(Demand)}$ (B.1)
$P_{mt}$ $=\theta_{0}+\tilde{\gamma}_{mt}Q_{mt}+\theta_{2}^{\prime}\mathbf{W}_{mt}+\theta_{3}I_{mt}+\vartheta_{mt}\quad\text{(Supply)}$ (B.2)
where $\mathbf{Z}_{mt}$ is a $d_{z}\times 1$ vector of demand shifters, $\mathbf{W}_{mt}$ is a $d_{w}\times 1$ vector of cost shifters, $I_{mt}$ is a scalar merger indicator, and $\varphi_{mt}$ and $\vartheta_{mt}$ represent unobserved demand and supply-side shocks, respectively. The parameters $\boldsymbol{\alpha}$ and $\boldsymbol{\theta}$ are $(2+d_{z})\times 1$ demand and supply parameter vectors, respectively.
@@
$\mathbb{E}[\mathbf{W}_{mt}(Q_{mt}-\alpha_{0}-\alpha_{1}P_{mt}-\alpha_{2}^{\prime}\mathbf{Z}_{mt})]=\mathbf{0}_{d_{w}\times 1}$ (B.8)
@@
$\mathbb{E}[\mathbf{Z}_{mt}(Q_{mt}-\alpha_{0}-\alpha_{1}P_{mt}-\alpha_{2}^{\prime}\mathbf{Z}_{mt})]=\mathbf{0}_{d_{z}\times 1}$ (B.9)
@@
$\mathbb{E}[\mathbf{Z}_{D,mt}\mathbf{X}_{D,mt}^{\prime}]\boldsymbol{\alpha}=\mathbb{E}[\mathbf{Z}_{D,mt}Q_{mt}]$ (B.14)
@@
$\mathbb{E}[\mathbf{Z}_{S,mt}(P_{mt}-\mathbf{X}_{S,mt}^{\prime}\boldsymbol{\theta})]=\mathbf{0}_{(1+d_{z}+1+d_{w})\times 1}$ (B.15)
@@
$\mathbb{E}[\mathbf{Z}_{S,mt}\mathbf{X}_{S,mt}^{\prime}]\boldsymbol{\theta}=\mathbb{E}[\mathbf{Z}_{S,mt}P_{mt}]$ (B.16)
@@
$\text{rank}(\mathbb{E}[\mathbf{Z}_{D,mt}\mathbf{X}_{D,mt}^{\prime}])=2+d_{z}$ (B.17)
@@
The order conditions for identification are as follows. For the demand equation, we have $1+d_{w}+d_{z}$ instruments and $2+d_{z}$ parameters, requiring $1+d_{w}+d_{z}\geq 2+d_{z}$, which simplifies to $d_{w}\geq 1$. For the supply equation, we have $2+d_{w}+d_{z}$ instruments and $3+d_{w}$ parameters, necessitating $2+d_{w}+d_{z}\geq 3+d_{q}$, which reduces to $d_{z}\geq 1$.
@@ C.2 Structural Estimates with Fixed Effects
The structural estimates using fixed effects without synthetic weights tell a dramatically different story from our preferred first-difference results. Standard GMM with fixed effects suggests that all three mergers generated significant pro-competitive effects: DL-NW and AA-US show large negative aggregate price effects (-\$34.2 and -\$41.9 per hundred dollars, respectively). At the same time, UA-CO demonstrates positive but significant effects (\$29.0). These results suggest that airline mergers uniformly benefited consumers through substantial price reductions.
@@
However, this finding changes completely when we use a synthetic GMM. Under synthetic GMM, the large, significant effects largely disappear, and the estimates become remarkably similar to our preferred first-difference results from the main text. Most coefficients lose statistical significance, and the magnitude of effects decreases substantially. For instance, the DL-NW aggregate price effect shrinks from -34.2 cents (highly significant) to 3.4 cents (insignificant), while the AA-US effect moderates from -41.9 cents to -21.5 cents.

### v6

@@ 5.3 Second Bridge: Quasi-Experimental to Structural
Our merger analysis faces a challenge: the number of treated markets relative to controls is small, making it difficult to obtain meaningful, precise parameter estimates. In our sample, for the DL-NW merger, there are only 14 unique treated markets compared to 822 control markets. The numbers for UA-CO and AA-US mergers are 6 and 768 markets, and 12 and 767 markets, respectively. This limited treatment variation may affect the estimates of efficiency gains and changes in conduct, as well as their precision.
Furthermore, selecting suitable control markets for merger analysis is challenging. The ideal control group should closely mirror the counterfactual evolution of treated markets in the absence of the merger, thereby satisfying the parallel trends assumption that is crucial for DiD estimation. As Ashenfelter et al. (2009) observe, researchers face an inherent tradeoff: control markets must be geographically distant enough to avoid merger spillovers yet similar enough to face comparable demand and cost conditions. This tradeoff presents the “similarity-independence tension” (Ashenfelter et al., 2009) in control market selection.
Our use of first-difference methods compounds the estimation challenge. Griliches and Mairesse (1999) demonstrate that first-difference estimators, while addressing fixed effects, can substantially reduce identifying variation. Our results confirm this pattern: first-difference specifications consistently yield larger standard errors and less significant coefficients compared to fixed effects approaches (Table C.2). However, fixed effects estimates may be biased because of the unobserved market characteristics correlated with the merger treatment.
We propose a solution that, in keeping with the paper’s overarching goal, bridges the quasi-experimental literature with structural econometrics. Our approach, which we term synthetic GMM, combines attractive features of synthetic control methods (Abadie and Gardeazabal, 2003) and synthetic DiD (Arkhangelsky et al., 2021) with structural estimation. The key innovation is straightforward and tractable. Like synthetic DiD, our method constructs the market weights to match pre-treatment trends of control markets with those for the treated markets, and time weights to balance pre- and post-merger periods for control markets. Then we use these synthetic DiD weights to reweight each market’s contribution to the moment conditions.
@@
3:Construct residualized pre-merger prices $\bar{P}_{tr,pre}$ (treated averages), and $\bar{P}_{co,pre}$ (control matrix) and post-merger prices $\bar{P}_{co,post}$
4:Solve for market weights $\hat{\omega}$ and time weights $\hat{\tau}$ following Arkhangelsky et al. (2021)
5:Define observation weights: $\omega_{mt}=\hat{\omega}_{m}$ if $m\in\mathcal{C}$, else 11; $r_{mt}=\hat{\tau}_{t}$ if pre-treatment, else 11
@@
6:Synthetic moments: $\tilde{g}(\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{\lambda})=\frac{1}{MT}\sum_{i,t}\omega_{mt}\cdot r_{mt}\cdot g(P_{mt},\boldsymbol{W}_{mt},\boldsymbol{Z}_{mt},Q_{mt};(\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{\lambda}))$
7:Synthetic GMM: $(\hat{\boldsymbol{\theta}},\hat{\boldsymbol{\alpha}},\hat{\boldsymbol{\lambda}}):=\arg\min_{(\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{\lambda})}\tilde{g}((\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{\lambda}))^{\top}\hat{W}\tilde{g}((\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{\lambda}))$
8:Bootstrap Inference:
9:for $b=1,\ldots,B$ do
@@
10:  Resample units with replacement (preserving time series and treatment status)
11:  Estimate weights $\hat{\omega}^{(b)},\hat{\tau}^{(b)}$ for bootstrap sample
@@
12:  Estimate synthetic GMM $(\hat{\boldsymbol{\theta}},\hat{\boldsymbol{\alpha}},\hat{\boldsymbol{\lambda}})^{(b)}$ using bootstrap weights
13:end for
@@
14:Compute bootstrap standard errors from $\{(\hat{\boldsymbol{\theta}},\hat{\boldsymbol{\alpha}},\hat{\boldsymbol{\lambda}})^{(b)}\}_{b=1}^{B}$
@@
More formally, our synthetic GMM approach proceeds in two main steps, as shown in the Table 6. First, following Arkhangelsky et al. (2021), we construct market-specific weights $\hat{\omega}$ that align pre-merger price trends between control and treated markets, and time-specific weights $\hat{\tau}$ that balance pre-merger periods. Like synthetic control methods, we find weights such that the weighted average of control market residualized prices closely matches treated market residualized prices: $\sum_{m\in\text{control}}\hat{\omega}_{m}P_{mt}\approx\frac{\sum_{m\in\text{treated}}P_{mt}}{N_{\text{treated}}}$ across all pre-merger periods $t$. The time weights similarly balance temporal variation for the control markets.
@@
Second, we integrate these weights directly into our GMM estimation by reweighting each moment condition, which we refer to as synthetic GMM. Instead of the standard sample moments $\frac{1}{MT}\sum_{m,t}g(\cdot;\cdot)$, we compute synthetic-weighted moments $\frac{1}{MT}\sum_{m,t}\omega_{mt}r_{mt}g(\cdot;\cdot)$, where $\omega_{mt}$ applies market weights to control units and $r_{mt}$ applies time weights to pre-treatment periods. This reweighting ensures that parameter identification emphasizes the most informative control observations while preserving the structural interpretation of synthetic GMM estimates. For inference, under the assumption that these weights are consistent, we implement block bootstrap resampling to jointly estimate the weights and structural parameters, accounting for uncertainty in both components.
@@
Table 7 displays the estimates. The estimates suggest heterogeneity in merger effects across the three airline consolidations, with notably greater precision than those in Table 5. The weighting approach yields smaller standard errors across all parameters, while maintaining the same qualitative conclusions about merger impacts. The estimates for the UA-CO merger suggest increased strategic conduct among remaining competitors by 0.049 (p-value 0.1), suggesting enhanced coordination in the post-merger market. Simultaneously, the merger appears to have reduced efficiency, i.e., increased cost, by \$13.5, which represents approximately 5.9% of the average pre-treatment price of \$228.25.
@@
These findings underscore the importance of rigorous merger analysis using a structural approach. The synthetic GMM approach can improve our ability to detect changes in conduct that might be missed with standard GMM. The results suggest that airline mergers were not pro-competitive as they generated limited efficiency gains while facilitating coordination.
@@ 6 Conclusion
This paper establishes a connection between quasi-experimental and structural approaches to merger evaluation. We build a two-way methodological bridge. The first bridge shows that the widely used DiD estimation equation emerges as the reduced form of a structural demand and supply equation, which, in turn, clarifies why DiD identification may fail in the merger context. The second bridge demonstrates that we can leverage recent advances in quasi-experimental methods, particularly synthetic control methods, to improve the estimation of structural models (synthetic GMM). This bridge reveals that these seemingly distinct methodologies represent different levels of economic detail applied to the same underlying relationships, and offers a unified framework that combines the interpretability of DiD with the robustness of structural identification.
We demonstrate that the DiD coefficient is a composite parameter that conflates cost efficiencies, economies of scale, and strategic conduct changes—effects that can only be separated through structural decomposition. Consequently, mergers pose fundamental identification challenges for DiD approaches: when mergers alter strategic behavior, the error structure in DiD specifications becomes endogenous to treatment status, leading to systematic bias. Our structural framework addresses these limitations while enabling researchers to construct DiD-equivalent measures from structural parameters, obtaining both methodological robustness and economic interpretability.
@@
Although focused on airlines, our bridging approach is broadly applicable to analyzing mergers in concentrated industries where strategic conduct plays a central role. The methodological framework provides a foundation for robust merger retrospectives that can inform both academic research and antitrust policy (Asker and Nocke, 2021).
@@ B Proofs of the Identification Results
$Q_{mt}$ $=\alpha_{0}+\alpha_{1}P_{mt}+\alpha_{2}^{\top}\mathbf{Z}_{mt}+\varphi_{mt}\quad\text{(Demand)}$ (B.1)
$P_{mt}$ $=\theta_{0}+\tilde{\gamma}_{mt}Q_{mt}+\theta_{2}^{\top}\mathbf{W}_{mt}+\theta_{3}I_{mt}+\vartheta_{mt}\quad\text{(Supply)}$ (B.2)
where $\mathbf{Z}_{mt}$ is a $d_{z}\times 1$ vector of demand shifters, $\mathbf{W}_{mt}$ is a $d_{w}\times 1$ vector of cost shifters, $I_{mt}$ is a scalar merger indicator, and $\varphi_{mt}$ and $\vartheta_{mt}$ represent unobserved demand and supply-side shocks, respectively. The parameters $\boldsymbol{\alpha}=(\alpha_{0},\alpha_{1},\alpha_{2}^{\top})$ and, with some abuse of notations, $\boldsymbol{\theta}=(\theta_{0},\tilde{\gamma}_{mt},\theta_{2}^{\top},\theta_{3})$ are $(2+d_{z})\times 1$ and $(3+d_{w})\times 1$ demand and supply parameter vectors, respectively.
@@
$\mathbb{E}[\mathbf{W}_{mt}(Q_{mt}-\alpha_{0}-\alpha_{1}P_{mt}-\alpha_{2}^{\top}\mathbf{Z}_{mt})]=\mathbf{0}_{d_{w}\times 1}$ (B.8)
@@
$\mathbb{E}[\mathbf{Z}_{mt}(Q_{mt}-\alpha_{0}-\alpha_{1}P_{mt}-\alpha_{2}^{\top}\mathbf{Z}_{mt})]=\mathbf{0}_{d_{z}\times 1}$ (B.9)
@@
$\mathbb{E}[\mathbf{Z}_{D,mt}\mathbf{X}_{D,mt}^{\top}]\boldsymbol{\alpha}=\mathbb{E}[\mathbf{Z}_{D,mt}Q_{mt}]$ (B.14)
@@
$\mathbb{E}[\mathbf{Z}_{S,mt}(P_{mt}-\mathbf{X}_{S,mt}^{\top}\boldsymbol{\theta})]=\mathbf{0}_{(1+d_{z}+1+d_{w})\times 1}$ (B.15)
@@
$\mathbb{E}[\mathbf{Z}_{S,mt}\mathbf{X}_{S,mt}^{\top}]\boldsymbol{\theta}=\mathbb{E}[\mathbf{Z}_{S,mt}P_{mt}]$ (B.16)
@@
$\text{rank}(\mathbb{E}[\mathbf{Z}_{D,mt}\mathbf{X}_{D,mt}^{\top}])=2+d_{z}$ (B.17)
@@
The order conditions for identification are as follows. For the demand equation, we have $1+d_{w}+d_{z}$ instruments and $2+d_{z}$ parameters, requiring $1+d_{w}+d_{z}\geq 2+d_{z}$, which simplifies to $d_{w}\geq 1$. For the supply equation, we have $2+d_{w}+d_{z}$ instruments and $3+d_{w}$ parameters, necessitating $2+d_{w}+d_{z}\geq 3+d_{w}$, which reduces to $d_{z}\geq 1$.
@@ C.2 Structural Estimates with Fixed Effects
The structural estimates using fixed effects without synthetic weights tell a dramatically different story from our preferred first-difference results. Standard GMM with fixed effects suggests that all three mergers generated significant pro-competitive effects: DL-NW and AA-US show large negative price effects (-\$34.2 and -\$41.9 per hundred dollars, respectively). At the same time, UA-CO demonstrates positive but significant effects (\$29.0). These results suggest that airline mergers uniformly benefited consumers through substantial price reductions.
@@
However, this finding changes completely when we use a synthetic GMM. Under synthetic GMM, the large, significant effects largely disappear, and the estimates become remarkably similar to our preferred first-difference results from the main text. Most coefficients lose statistical significance, and the magnitude of effects decreases substantially. For instance, the DL-NW price effect shrinks from -34.2 cents (highly significant) to 3.4 cents (insignificant), while the AA-US effect moderates from -41.9 cents to -21.5 cents.